### v3

@@ 2 Data and Methods
Psuedo-time assignment, also known as trajectory inference, is the ordering of Single-cell RNA Sequencing experiments into trajectories and steady states. We have developed a novel pesudo-time assignment algorithm that is based on our modeling methodology, and is analogous an expectation-maximization procedure. Initially, the reconstructed network is set to the model that contains all the candidate edges, with some initial logic. For example, if edge signs are available, a rule like ”inhibitors win over activators’ can be used. Next, using the initial network, network trajectories from the states corresponding to the observed Boolean states of the cells are generated: each trajectory is extended until an attractor is reached, and also includes all attractor states. Each Boolean cell state is then mapped to its closest state in the trajectories that were obtained. Contiguous sequences of states are assigned to the corresponding trajectories time points, where trajectories of length 1 are set as steady states. Using the new assignment of pseudo-time, a new network is reconstructed from the data. The process repeats itself until the value of the solution stops improving. In the EM analogy, the latent variables are the cell assignments to pseudo-time points, and the network edges and logic are the model parameters.
@@ 3 Results
Next, we applied our pseudo-time inference algorithm to the HSC network of Bonzanni et al. (Bonzanni et al., 2013), using GEO dataset GSE75478 which profiles human hematopoeitic stem cells in early differentiation, in order to try to reconstruct the network from single-cell data. The raw data from the file GSE75478_transcriptomics_raw_filtered_I2.csv.gz was preprocessed using the R package scuttle, followed by library-size normalization and log transformation. Binarized values were obtained using the R Binarize package and the kMeans method, with default arguments. As an initial network for the pseudo-time inference algorithm, we used the edge signs published in Bonzanni et al. with the ”inhibitors win” rule for combinatorial interactions. The initial rate of mismatches was 0.17, and after convergence it was reduced to 0.09. All the inferred interactions matched the signs of the edges published in Bonzanni et al., though the number of inferred interactions was smaller, likely due to the noisy nature of scRNA-Seq data. We conclude from both experiment that accurate network reconstruction is possible from varying data types and in the presence of asynchronous dynamics.

### v4

@@ 2 Data and Methods
Psuedo-time assignment, also known as trajectory inference, is the ordering of Single-cell RNA Sequencing experiments into trajectories and steady states. We have developed a novel pesudo-time assignment algorithm that is based on our modeling methodology, and is analogous an expectation-maximization procedure. Initially, the reconstructed network is set to the model that contains all the candidate edges, with some initial logic. For example, if edge signs are available, a rule like ”inhibitors win over activators’ can be used. Next, using the initial network, network trajectories from the states corresponding to the observed Boolean states of the cells are generated: each trajectory is extended until an attractor is reached, and also includes all attractor states. Each observed Boolean cell state is then mapped to its closest state in the trajectories that were obtained. Contiguous sequences of states are assigned to the corresponding trajectories time points, where trajectories of length 1 are set as steady states. Using the assignment of pseudo-time, a new network is reconstructed from the data, and a set of de-noised cell states is obtained, from which a new pseudo-time is generated. Using the new pseudo-time, a network is fitted to the data. The process repeats itself until the value of the solution stops improving. In the EM analogy, the latent variables are the cell assignments to pseudo-time points, and the network edges and logic are the model parameters.
@@ 3 Results
Next, we applied our pseudo-time inference algorithm to the HSC network of Bonzanni et al. (Bonzanni et al., 2013), using GEO dataset GSE75478 which profiles human hematopoeitic stem cells in early differentiation, in order to try to reconstruct the network from single-cell data. The raw data from the file GSE75478_transcriptomics_raw_filtered_I2.csv.gz was preprocessed using the R package scuttle, followed by library-size normalization and log transformation. Binarized values were obtained using the R Binarize package and the kMeans method, with default arguments. As an initial network for the pseudo-time inference algorithm, we used the edge signs published in Bonzanni et al. with the ”inhibitors win” rule for combinatorial interactions. The initial rate of mismatches was 0.158, and after convergence it was reduced to 0.126. All the inferred interactions matched the signs of the edges published in Bonzanni et al., though for the regulation of RUNX1, the activiating effect of its regulators seem to require two active regulators (Table 2). We conclude from both experiment that accurate network reconstruction is possible from varying data types and in the presence of asynchronous dynamics.
@@
[…]
| GATA1 | SPI1 | repression | repression |
| FLI1 | SPI1 | activation | activation |
| RUNX | SPI1 | activation | activation |
| SPI1 | GATA1 | repression | repression |
| TAL1 | GATA1 | activation | activation |
| GATA1 | TAL1 | activation | activation |
| ERG1 | TAL1 | activation | activation |
| GATA2 | TAL1 | activation | activation |
| GATA1 | FLI1 | repression | repression |
| ERG1 | FLI1 | activation | activation |
| TAL1 | FLI1 | activation | activation |
| RUNX1 | ERG | activation | activation |
| TAL1 | ERG | repression | repression |
| TAL1 | CBFA2T3 | activation | activation |
| SPI1 | RUNX1 | activation | activation |
| TAL1 | RUNX1 | activation | activation |
| GATA2 | RUNX1 | activation | activation |
[…]